what is the total cost of a pair of jeans that are priced $40.00 if there is a 6.5% sales tax?
solve for x: 4(3x-5)=7(2x+3)


Answer by rfer(16322)

A)
40*1.065=42.60
B)
12x-20=14x+21
add 20
12x=14x+21+20
subtract 14x
12x-14x=21+20
-2x=41
x=-20.5